OpenStudy (anonymous):

If we toss one die, what is the probability that it is one of the numbers 1, 2, 3, 4, 5, or 6? (Points : 4) 1/6 5/6 1 0

OpenStudy (anonymous):

P(1 or 2 or 3 or ... or 6) = 1/6 + 1/6 + 1/6 + 1/6 + 1/6 + 1/6 = 6/6 = 1